# QUITE: A Query Rewrite System Beyond Rules with LLM Agents

### v2

@@ 4. LLM Agent-Based Query Rewriter
Figure 4. Reasoning Chain of MDP-based Reasoning Agent
@@ 4.1. MDP-based Reasoning Agent
Query rewrite is fundamentally a sequential decision problem (Finance and Gardarin, 1991b; Zhou et al., 2021), each transformation step changes the query structure and limits future optimization options. Reasoning LLM agents (e.g., DeepSeek-R1 (Guo et al., 2025)), capable of chain-of-thought planning and multi-step logical inference, appear well-suited to this task. However, our experiments show that they still produce 13.0% nonequivalent outputs and yield minimal execution improvements (Section 7.2). We observe that query rewrite mirrors a Markov Decision Process (MDP), where states are query forms, actions are rewrite steps, and rewards are cost reductions (Puterman, 1990; Liu et al., 2023; Wang, 2025). By framing rewrite in this way, we provide the reasoning LLM agent with a clear, stepwise optimization goal that guides each inference step. Therefore, we design an MDP-based reasoning agent to produce high-performing rewrites.
We model the query rewrite process as a Markov Decision Process $(\mathcal{S},\mathcal{A},\mathcal{T},r,\gamma)$. Starting from the original query $Q$, each state $S_{t}\in\mathcal{S}$ represents the current form of the query with any semantical-level rewrite attempts. At each timestep $t$, the agent chooses either a refinement action $R_{t}$ (e.g., join reordering, predicate pushdown) or a terminal action $\mathcal{A}$ that emits the final rewritten query. This yields the deterministic transition:
@@
To guide the search toward efficient rewritten queries, we define a domain‐specific reward function once $S_{t}$ transitions to $S_{t+1}$ under $R_{t}$:
@@
As illustrated in Figure 4, the agent transitions from the initial state $S_{0}$ through a sequence of refinement actions $\{R_{0},R_{1},\dots\}$, and finally applies the terminal action $A$ to emit the final rewritten query. Concretely, at each timestep $t$, the agent is prompted with the current SQL form $S_{t}$ and analyzes the query structure via its internal chain‐of‐thought, scores candidate refinements according to the expected cost reduction, and samples the highest‐scoring $R_{t}$.
@@ 4.2. FSM-based Query Rewrite
Single-agent approaches often produce suboptimal or invalid rewrites because a single agent is responsible for managing the entire rewrite pipeline (e.g., propose rewrites, verify validity, and assess efficiency), which overloads the model’s context capacity and ultimately increases the risk of hallucination. Inspired by multi-agent frameworks that assign specialized agents to distinct tasks (Hong et al., 2023), we break down query rewrite into discrete stages, each is handled by a dedicated LLM agent. To ensure the multi-agent system’s stability and correctness (Han et al., 2024), we implement a Finite State Machine (FSM) to orchestrate stable stage transitions.
@@ 4.2.1. Query Rewrite Process Decomposition
  • $\mathcal{M}:\mathcal{Q}\times\mathcal{E}\rightarrow\mathcal{S}$ is a transmission that breaks down the input query $Q\in\mathcal{Q}$ into intermediate states $s\in\mathcal{S}$.
@@
4.2.2. Specialized LLM Agents
As shown in Table 1, we deploy three specialized LLM agents alongside the core MDP-based Reasoning Agent in Section 4.1 to drive the rewrite pipeline for defined FSM stages. Then we propose an LLM agent-based query rewrite methodology detailed in Algorithm 1:
  • Decision stage ($S_{3}$): In this stage, the Decision Agent reassesses the rewritten SQL’s efficiency with comprehensive report, and determines whether the rewrite process should continue or terminate (Lines 13-19).
@@ 9             $\langle Q^{E},\text{EarlyStop}\rangle\leftarrow\text{Enhance-Query}(\mathcal{M},Q^{P},\mathcal{A})$;
10             $S\leftarrow\text{EarlyStop}\vspace{0.25pt}?\vspace{0.25pt}\textsc{Termination}:\textsc{Verification}$;
@@ 20             $\mathcal{R}\leftarrow\text{Generate-Report}(Q^{E},Q_{0},\mathcal{M},T)$;
[…]
@@ 29       $t\leftarrow t+1$;
Algorithm 1 LLM-Agent-based Query Rewrite Algorithm
Rewrite Agent. After the Reasoning agent generates a chain, the Rewrite agent extracts proposals from each node, groups them into four categories matching our structured knowledge base, evaluates each proposal’s expected reward, and selects the top SQL candidate. Since this candidate may introduce new structural complexities, it cannot be used directly for verification. Therefore, the rewrite agent undertakes an SQL enhancement process to uncover additional rewrite opportunities. The prompt template for this process is shown below:
[…]
1. Push down predicates or use a CTE structure
2. Simplify calculations and remove redundant predicates
3. Apply other efficiency-preserving transformations
If you think the SQL candidate is well-optimized, just return the original version as the rewritten SQL.
@@ 3. Apply other efficiency-preserving transformations
Assistant Agent. The Assistant agent uses the hybrid SQL corrector (Section 5.2) to ensure the syntactically correctness and semantically equivalence of queries. This is performed in two steps:
  • Syntax correction. If syntax errors occur, it attempts up to $K_{\max}$ repairs. If the query remains uncorrected after these attempts, the current FSM iteration is aborted, and control returns to the reasoning stage.
  • Equivalence correction. Once syntax error is corrected, the agent submits the query pair $\langle Q_{\text{orig}},Q_{\text{new}}\rangle$ to verify the equivalence and correct nonequivalent queries.
This two-phase process ensures that only valid, semantically equivalent rewrites advance to the Decision stage. Although equivalence verification may return the original format, it avoids the far greater cost of returning nonequivalent SQL while preventing performance degradation from excessive loop iterations.
@@
(i) Generate Comprehensive Report. The Decision agent analyzes the rewritten query across four dimensions, including cost changes, plan characteristics, resource utilization, and other improvements. This multi-dimensional approach overcomes the limitations of relying solely on EXPLAIN cost estimates, preventing the system from accepting misleading or suboptimal rewrites.
(2) Evaluation and Feedback Loop. Based on the generated report, the Decision agent gives a binary judge to decide whether the rewrite process should terminate or not. If True, the Decision agent terminates the FSM process; If False, it retrieves higher-confidence rewrite proposals from the knowledge base, then stores the current query, report, and retrieved knowledge in an agent shared buffer for the next FSM iteration. We deliberately avoid introducing knowledge base retrieval in the initial state, because reasoning models have input content limitations—excessive guidance information can reduce the generation freedom, potentially compromising the quality of reasoning proposal chain (Guo et al., 2025). Finally, the FSM enters the Termination stage and outputs the final rewritten query along with its comprehensive optimization report.
@@ 5.1. Structured Knowledge Base
Since DBAs encode numerous scenario-specific rewrite strategies that LLMs cannot learn from sparse examples and that fail to generalize across diverse query contexts, we consolidate these expert rules into a structured, scenario-indexed knowledge base. To this end, preparing the knowledge base involves the following steps: (i) collecting SQL rewrite knowledge from various resources, (ii) filtering untrustworthy content, (iii) enriching the entries, and (iv) retrieving the most relevant rules for each rewrite task. The Knowledge Base $\mathcal{K}$ is defined as follows:
Definition 1 (Knowledge Base). A Knowledge Base $\mathcal{K}$ is a set of question and answer (Q&A) pairs, denoted as $\mathcal{K}=\{\langle q_{1},a_{1}\rangle,\langle q_{2},a_{2}\rangle,\allowbreak...,\langle q_{n},a_{n}\rangle\}$. Each Q&A pair $\langle q_{i},a_{i}\rangle$ consists of:
  • A question $q_{i}=\langle Text_{que_{i}},SQL_{que_{i}}\rangle$, where $Text_{que_{i}}$ describes a specific SQL rewrite issue, and $SQL_{que_{i}}$ is the original SQL.
@@
Step 1: Collecting rewrite knowledge. Our knowledge is primarily sourced from official database documentation and the database community. Since rewrite strategies are often indirectly presented and scattered throughout official documents (pos, [n.d.]b; mys, [n.d.]; sql, [n.d.]; ora, [n.d.]), we employ LLMs to identify and summarize them into key points. We also collect knowledge from StackOverflow (sta, 2025), as it has abundant real-world query rewrite examples and insights accumulated over decades. We retrieve all Q&A units tagged with “query rewrite” or “query optimization” that include at least one answer. In total, we collected 3,432 valid Q&A units.
@@
Step 3: Enhancing the knowledge. An LLM first summarizes questions and user’s responses from website Q&A units. These summaries are then augmented with corroborating information from official documents. To do this, we embed user response summaries using Sentence Transformers (Reimers and Gurevych, 2019). These embeddings are then used to query a pre-indexed knowledge base of extracted points, retrieving the top three matches via cosine similarity. Finally, an LLM evaluates these matches to confirm that they effectively explain the rewrite strategy underlying the Q&A unit. Suitable matches are incorporated, enriching Q&A units with concise explanations that aid future retrieval.
@@
Step 4: Retrieving relevant knowledge. We classify these units into four categories: join optimization, constant folding, predicate simplification and others, which capture the essential aspects of query rewriting and provides a foundational structure for enabling precise retrieval of information. Since knowledge bases are often context-specific, we use the BM25 algorithm (Robertson et al., 2009) to retrieve the most relevant documents by computing the document relevance.
5.2. Hybrid SQL Corrector
Equivalence Error Correction. Verifying the equivalence between any pair of SQL queries is known to be an NP-hard challenge (Aho et al., 1979), and whether this problem admits a polynomial-time solution remains an open question. While numerous verification methods (Chu et al., 2017a; Wang et al., 2022, 2024; Ding et al., 2023) rooted in algebraic or symbolic reasoning have been proposed to prove equivalence, they typically operate on constrained query sets (e.g., specific operators, advanced SQL constructs). Recent studies have highlighted the potential of LLMs in deducing query equivalence with high confidence (Singh and Bedathur, 2024; Zhao et al., 2023). However, relying solely on LLMs carries inherent risks. Therefore, we integrate traditional verification tools and LLM capabilities in a two-stage process, forming a hybrid approach that achieves broader and more reliable equivalence judgments.
  • Tool-based Verification. We first use SQLSolver (Ding et al., 2023), an advanced verification tool, for an initial check. Given a pair of queries, $\langle Q_{orig},\allowbreak Q_{rewr}\rangle$, SQLSolver determines their equivalence as either equivalent, nonequivalent, or unknown (due to timeout or other limitations). For query pairs where SQLSolver’s outcome is unknown, verification proceeds to a subsequent stage, utilizing the proposed LLM-based approach.
  • LLM-based Verification and Correction. Leveraging their ability to understand semantic relationships, LLMs can effectively check SQL query equivalence and correct nonequivalent rewrites. However, directly applying LLMs faces challenges, as they may be unsure for some complex queries or incorrectly identify a pair of nonequivalent queries as equivalent. To mitigate these issues, we employ an iterative verification and correction process. In this process, the LLM repeatedly compares the original query $Q_{orig}$ with the rewrite candidate, systematically identifying discrepancies and applying refinements. This iterative loop continues until the LLM confidently establishes query equivalence or a predefined time budget is exhausted. If the time budget is reached without the LLM confidently establishing equivalence, the system defaults to returning the original query ($Q_{rewri}=Q_{orig}$), ensuring that system performance is not compromised by unresolved complex rewrites. Our experiments on diverse workloads show that our method achieves higher rewrite equivalence rates than baselines (Figure 6).
5.3. Agent Memory Buffer
Effective context sharing is crucial in LLM-agent based query rewrite workflows. Simply logging the entire prompt history is problematic: (1) Hallucinations due to redundant context. Lengthy SQL statements and redundant reasoning tokens can distract LLM, causing it to generate incorrect content (Shi et al., 2023). (2) Performance degradation under long-context overhead. Unbounded history growth incurs quadratic attention costs and increases inference latency (Jiang et al., 2023; Kuratov et al., 2024).
To address these, we propose an agent memory buffer $\mathcal{B}$, containing memory slices =$\mathcal{B}=\{m_{1},m_{2},...,m_{n}\}$, where each slice $m_{i}$ is dedicated to a specific category of critical information (e.g., query information, rewrite proposals, retrieval knowledge). By extracting only essential messages across the rewrite process, our buffer curbs hallucinations and bounds context growth, thereby improving both correctness and efficiency in single-agent SQL rewrite workflows.
@@ 5.3. Agent Memory Buffer
NOT MATERIALIZE /*+ NOT MATERIALIZE(table) */ Inlines the CTE to avoid materialization overhead.

### v3

@@ 4.1. MDP-based Reasoning Agent
Query rewrite is fundamentally a sequential decision problem (131472; lr), each transformation step changes the query structure and limits future optimization options. Reasoning LLM agents (e.g., DeepSeek-R1 (DeepSeek-r1)), capable of chain-of-thought planning and multi-step logical inference, appear well-suited to this task. However, our experiments show that they still produce 13.0% nonequivalent outputs and yield minimal execution improvements (Section 7.2). We observe that query rewrite mirrors a Markov Decision Process (MDP), where states are query forms, actions are rewrite steps, and rewards are cost reductions (mdp; liu2023reason; tutorial). By framing rewrite in this way, we provide the reasoning LLM agent with a clear, stepwise optimization goal that guides each inference step. Therefore, we design an MDP-based reasoning agent to produce high-performing rewrites.
We model the query rewrite process as a Markov Decision Process $(\mathcal{S},\mathcal{A},\mathcal{T},r,\gamma)$. Starting from the original query $Q$, each state $S_{t}\in\mathcal{S}$ represents the current form of the query with any semantic-level rewrite attempts. At each timestep $t$, the agent chooses either a refinement action $R_{t}$ (e.g., join reordering, predicate pushdown) or a terminal action $\mathcal{A}$ that emits the final rewritten query. This yields the deterministic transition:
@@
To guide the search toward efficient rewritten queries, we define a domain‐specific reward function once $S_{t}$ transitions to $S_{t+1}$ under a refinement action $R_{t}$:
@@
As shown in Figure 4, the agent transitions from the initial state $S_{0}$ through a sequence of refinement actions $\{R_{0},R_{1},\dots\}$, and finally applies the terminal action $A$ to emit the final rewritten query. Concretely, at each timestep $t$, the agent is prompted with the current SQL form $S_{t}$ and analyzes the query structure via its internal chain‐of‐thought, scores candidate refinements based on the expected cost reduction, and samples the highest‐scoring $R_{t}$.
@@ 4.2. FSM-based Query Rewrite
Single-agent approaches often produce suboptimal or invalid rewrites because a single agent is responsible for managing the entire rewrite pipeline (e.g., propose rewrites, verify validity, and assess efficiency), which overloads the model’s context capacity and ultimately increases the risk of hallucination. Inspired by multi-agent frameworks that assign specialized agents to distinct tasks (metagpt), we break down query rewrite into discrete stages, each of which is handled by a dedicated LLM agent. To ensure the multi-agent system’s stability and correctness (multi-agent-chanllenges), we implement a Finite State Machine (FSM) to orchestrate stable stage transitions.
@@ 4.2.1. Query Rewrite Process Decomposition
  • $\mathcal{M}:\mathcal{Q}\times\mathcal{E}\rightarrow\mathcal{S}$ is a transition that breaks down the input query $Q\in\mathcal{Q}$ into intermediate states $s\in\mathcal{S}$.
@@
4.2.2. Specialized LLM Agents
As shown in Table 1, we deploy three specialized LLM agents alongside the core MDP-based Reasoning Agent in Section 4.1 to drive the rewrite pipeline for the defined FSM stages. Then we propose an LLM agent-based query rewrite methodology detailed in Algorithm 1:
  • Decision stage ($S_{3}$): In this stage, the Decision Agent reassesses the rewritten SQL’s efficiency with a comprehensive report and determines whether the rewrite process should continue or terminate (Lines 13-19).
@@ 9    $\langle Q^{E},\text{EarlyStop}\rangle\leftarrow\text{Enhance-Query}(\mathcal{M},Q^{P},\mathcal{A})$;
10    $S\leftarrow\text{EarlyStop}\vskip 0.25pt?\vskip 0.25pt\textsc{Termination}:\textsc{Verification}$;
@@ 20    $\mathcal{R}\leftarrow\text{Generate-Report}(Q^{E},Q_{0},\mathcal{M},T)$;
21    if $\text{Cost-Check}(Q^{E},\mathcal{R},\mathcal{M})$ then
@@ 29 $t\leftarrow t+1$;
33
Algorithm 1 LLM-based Query Rewrite Algorithm
Rewrite Agent. After the reasoning agent generates a chain, the rewrite agent extracts proposals from each node, groups them into four categories matching our structured knowledge base, evaluates each proposal’s expected reward, and selects the top SQL candidate. Since this candidate may introduce new structural complexities, it cannot be used directly for verification. Therefore, the rewrite agent undertakes an SQL enhancement process to uncover additional rewrite opportunities.
@@
Assistant Agent. The assistant agent uses the hybrid SQL corrector (Section 5.2) to ensure the syntactic correctness and semantically equivalence of queries. This is performed in two steps:
  • Syntax correction. If syntax errors occur, it attempts up to $K_{\max}$ repairs. If correction still fails, the current FSM iteration is aborted and control returns to the reasoning stage.
  • Equivalence correction. Once the syntax error is corrected, the agent submits the query pair $\langle Q_{\text{orig}},Q_{\text{new}}\rangle$ to verify the equivalence and correct nonequivalent queries.
This two-phase process ensures that only valid, semantically equivalent rewrites advance to the Decision stage. Although equivalence verification may return the original input SQL, it avoids the far greater cost of returning nonequivalent SQL while preventing performance degradation from excessive loop iterations.
@@
(i) Generate Comprehensive Report. The decision agent analyzes the rewritten query across four dimensions, including cost changes, plan characteristics, resource utilization, and other improvements. This multi-dimensional approach overcomes the limitations of relying solely on EXPLAIN cost estimates, preventing from accepting misleading or suboptimal rewrites.
(2) Evaluation and Feedback Loop. Based on the generated report, the decision agent makes a binary decision to decide whether the rewrite process should terminate or not. If True, the Decision agent terminates the FSM process; if False, it retrieves higher-confidence rewrite proposals from the knowledge base, then stores the current query, report, and retrieved knowledge in an agent shared buffer for the next FSM iteration. We deliberately avoid introducing knowledge base retrieval in the initial state, because reasoning models have input content limitations—excessive guidance information can reduce the generation freedom, potentially compromising the quality of reasoning proposal chain (DeepSeek-r1). Finally, the FSM enters the Termination stage and outputs the final rewritten query along with its comprehensive optimization report.
@@ 5.1. Structured Knowledge Base
Many query rewrite strategies developed by DBAs cannot be effectively utilized by LLMs. This limitation arises because such knowledge is sparsely represented in the LLMs’ training data, making it difficult for them to extract and apply these strategies without explicit guidance provided in the prompt. To address this, we construct a structured knowledge base to systematically manage various rewrite strategies. The construction process consists of four steps: (i) collecting SQL rewrite knowledge from diverse sources, (ii) filtering unreliable content, (iii) enriching the collected entries, and (iv) retrieving the most relevant rules for each rewrite. The Knowledge Base $\mathcal{K}$ is defined as follows:
Definition 1 (Knowledge Base). A Knowledge Base $\mathcal{K}$ is a set of question and answer (Q&A) pairs, denoted as $\mathcal{K}=\{\langle q_{1},a_{1}\rangle,\langle q_{2},a_{2}\rangle,\allowbreak...,\langle q_{n},a_{n}\rangle\}$. Each Q&A pair $\langle q_{i},a_{i}\rangle$ includes:
  • A question $q_{i}=\langle Text_{que_{i}},SQL_{que_{i}}\rangle$, where $Text_{que_{i}}$ describes a specific SQL rewrite issue in natural language, and $SQL_{que_{i}}$ is the original SQL query that the system needs to interpret or optimize.
@@
Step 1: Collecting rewrite knowledge. Our knowledge is primarily sourced from official database documentation and the database community. Since rewrite strategies are often indirectly presented and scattered throughout official documents (postgresqldocs; mysql_docs; sqlserver_docs; oracle_db_docs), we employ LLMs to identify and summarize them into key points. We also collect knowledge from Stack Overflow (stackoverflow), as it has abundant real-world query rewrite examples and insights accumulated over decades. We retrieve all Q & A units tagged with “query rewrite” or “query optimization” that include at least one answer. In total, we collected 3,432 valid Q & A units.
@@
Step 3: Enhancing the knowledge. An LLM first summarizes questions and user’s responses from website Q&A units. These summaries are then augmented with corroborating information from official documents. To do this, we embed user response summaries using Sentence Transformers (sentence-transformer). These embeddings are then used to query a pre-indexed knowledge base of extracted points, retrieving the top three matches via cosine similarity. Finally, an LLM evaluates these matches to confirm that they effectively explain the rewrite strategy underlying the Q&A unit. Suitable matches are incorporated, enriching Q&A units with concise explanations that aid future retrieval.
Table 2. Distribution of Knowledge UNits Across Categories
Join Constant Predicate CTE Others
122 2 21 20 76
@@
Step 4: Retrieving relevant knowledge. We classify these 241 high-quality units into five categories: predicate, constant, join, CTE and others, which capture the essential aspects of query rewrite and provide a foundation for precise retrieval (Table 2).s Since knowledge is often context-specific (i.e., most rewrite strategies depend heavily on factors such as schema definitions, data distribution and available indexes), its effectiveness may diminish under changing conditions. In contrast, context-independent knowledge is relatively rare and involves general-purpose optimizations such as predicate simplification. To identify the most relevant knowledge, we use the BM25 algorithm (bm25) to retrieve the most relevant documents by computing the document relevance.
@@
NOT_MATERIALIZE /*+ NOT_MATERIALIZE(table) */ Inlines the CTE to avoid materialization overhead.
5.2. Hybrid SQL Corrector
Equivalence Error Correction. Verifying the equivalence between any pair of SQL queries is known to be an NP-hard challenge (eq_np), and whether this problem admits a polynomial-time solution remains an open question. While numerous verification methods (cosette; wetune; qed; sqlsolver) rooted in algebraic or symbolic reasoning have been proposed to prove equivalence, they typically operate on constrained query sets (e.g., specific operators, advanced SQL constructs). Recent studies have highlighted the potential of LLMs in deducing query equivalence with high confidence (singh2024exploring; LLMSQLSolverCL). However, relying solely on LLMs carries inherent risks. Therefore, we integrate traditional verification tools and LLM capabilities in a two-stage process, forming a hybrid approach that achieves broader and more reliable equivalence judgments.
  • Tool-based Verification. We first use SQLSolver (sqlsolver), an advanced verification tool, for an initial check. Given a pair of queries, $\langle Q_{orig},\allowbreak Q_{rewr}\rangle$, SQLSolver determines their equivalence as either equivalent, nonequivalent, or unknown (due to timeout or other limitations). For query pairs where SQLSolver’s outcome is unknown, verification proceeds to a subsequent stage, utilizing the proposed LLM-based approach.
  • LLM-based Verification and Correction. Leveraging their ability to understand semantic relationships, LLMs can effectively check SQL query equivalence and correct nonequivalent rewrites. However, directly applying LLMs faces challenges, as they may be unsure for some complex queries or incorrectly identify a pair of nonequivalent queries as equivalent. To mitigate these issues, we employ an iterative verification and correction process. In this process, the LLM repeatedly compares the original query $Q_{orig}$ with the rewrite candidate, systematically identifying discrepancies and applying refinements. This iterative loop continues until the LLM confidently establishes query equivalence or a predefined time budget is exhausted. If the time budget is reached without the LLM confidently establishing equivalence, the system defaults to returning the original query ($Q_{rewr}=Q_{orig}$), ensuring that system performance is not compromised by unresolved complex rewrites. Our experiments on diverse workloads show that our method achieves higher rewrite equivalence rates than baselines (Figure 6).
5.3. Agent Memory Buffer
Effective context sharing is crucial in LLM-based workflows. Simply logging the entire prompt history is problematic: (1) Hallucinations due to redundant context. Lengthy SQL statements and redundant reasoning tokens can distract LLM, causing it to generate incorrect content (shi2023large). (2) Performance degradation under long-context overhead. Unbounded history growth incurs quadratic attention costs and increases inference latency (jiang2023longllmlingua; kuratov2024babilong).
To address these, we propose an agent memory buffer $\mathcal{B}$, containing memory slices $\mathcal{B}=\{m_{1},m_{2},...,m_{n}\}$, where each slice $m_{i}$ is dedicated to a specific category of critical information (e.g., query information, rewrite proposals). By extracting only essential messages across the rewrite process, our buffer curbs hallucinations and bounds context growth, thereby improving both correctness and efficiency in single-agent SQL rewrite workflows.